Can you check my math?
simply put, this is a mathematical setup of the 73% chance to win at first tower taken or consequently, the other team only has 27% chance to win average
the winning team has a potential chance to win of 100% , but just by existing, this brings the lowest chance to win at 46% for the winning team
let's look at the other side, however. The losing side has a 27% average, but, because it is a question of yes or no whether they win or not, the chance for victory is highest at 50%, making the match equal between the two teams
with 50% as the marker, that means that the losing team can have a minimum of 4% chance to win and a maximum of 50% chance to win
using this data and by player comparison, let's tally it from one extreme to the next
the range goes from 5 uncoordinated and unskilled players in their current division for the minimum values
to 5 lucky coordinated and skilled players that happened to be put into the same group
the values are as follows and I will put them in X / Y formats for uncoordinated to extremely good
5 / 0 = 4% chance of victory 4/ 1 = 13.2% chance of victory 3 / 2 = 22.4% chance of victory 2/3 = 31.6% chance of victory 1/4 = 40.8% chance of victory 0/5 = 50% chance of victory
then we can factor values in-between. Say a neutral player who does not lose but does not succeed either
I could make a complete list of probabilities, but i'm going to assume that there are no "unskilled" players for this set of numbers
the format is X = neutral / y = skilled
5 / 0 = 27% chance of victory 4 / 1 = 31.6% chance of victory 3/ 2 = 36.2% chance of victory 2/ 3 = 40.8% chance of victory 1 / 4 = 45.4% chance of victory 0 / 5 = 50% chance of victory
the numbers are there, but as well all know, it's usually going to be at least one unqualified player
so, i'll take the math for 1 unqualified and 4 neutral
and the match for 1 unqualified, 3 neutral and 1 skilled to show a comparison
the format is the same now with x / y / z
1 / 4 / 0 = 22.4% chance of victory
1/ 3 / 1 = 27% chance of victory, interestingly in a perfect world, the good and bad player cancel each other out, but gold leads do not do so as this is just statistical probability
now for the fun of it, let's say there's 2 bad, 2 neutral and 1 good (as this sometimes happens)
it drops the chance back down to 22.4 %
these numbers are quite terrifying when the team that has the objective has a 46% chance of victory just by existing...(as in lowest tier possible skill to win the match against the opponent)
and if all of the opponents are at the highest skill level, that chance becomes 100%
the simple formula is
[(a + b + c + d + e ) divided by ( 5 times ( highest possible win percentage))] times ( highest possible win percentage number)
where a through e are represented by player skill levels unskilled equal minimum win percentage to neutral win percentage to highest possible win percentage
any comments?
quotes:
"Stop that! You will damage my turrets!" Eureka!! "This behaviour is highly inefficient!" "By my calculations... " "I theorize... your defeat!" "I've applied the formula for success!" "They'll have a few screws loose by the end of this!" "Just as I hypothesized!"
"Fascinating, isn't it?" "Not enough variables... hmm... not nearly enough variables." "Every problem solved is another revealed."
"One step closer to greater understanding!"
"Back to the drawing board!" "I could do that better."